OpenStudy (anonymous):

What percent of 165 is 75

OpenStudy (phi):

what % of 165 = 75 write that as x*165 = 75 and we want to find "x" divide both sides by 165 \[ x\cdot \frac{165}{165}= \frac{75}{165} \\ x= 0.4545 \] now change 0.4545 to a percent (by multiplying it by 100)